OpenStudy (spring98):

The incorrect work of a student to solve an equation 2(y + 8) = 4y is shown below: Step 1: 2(y + 8) = 4y Step 2: 2y + 10 = 4y Step 3: 2y = 10 Step 4: y = 5 Which of the following explains how to correct Step 2 and shows the correct value of y? A. The equation should be y + 8 = 4y after division by 2; y = 4 B. The equation should be y + 8 = 4y after division by 2; y = 6 C. 2 should be distributed as 2y + 16; y = 4 D. 2 should be distributed as 2y + 16; y = 8

OpenStudy (anonymous):

HEYYYY

OpenStudy (spring98):

hi

OpenStudy (spring98):

can you help?

OpenStudy (anonymous):

can you go on gmail..

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

The key to this is the distributive property, which says that 2(x+y) = 2x+2y

OpenStudy (spring98):

i can't! today is the last day of school for me and i need to finish my assignments!

OpenStudy (anonymous):

Just use the distributive property to distribute the 2 for (y+8)

OpenStudy (spring98):

how?

OpenStudy (anonymous):

If 2(x+y)=2x+2y then 2(y+8)=what?

OpenStudy (spring98):

8?

OpenStudy (spring98):

is it 8?

OpenStudy (anonymous):

Just plug in "y" for x and "8" for y

OpenStudy (spring98):

it equals 4

OpenStudy (anonymous):

Can you show me HOW you're getting these numbers? I can't tell what you're doing wrong if you don't give me any work.

OpenStudy (spring98):

2x+2y=4

OpenStudy (anonymous):

The 2x+2y is just an example of how the distributive property works. It's completely unrelated to the problem.

OpenStudy (spring98):

16y?

OpenStudy (anonymous):

Don't add the y and the 8.

OpenStudy (spring98):

2

OpenStudy (spring98):

so you get two?

OpenStudy (anonymous):

2(y+8)=2(y)+2(8)

OpenStudy (spring98):

ho do i find the y? i need the why not the x!

OpenStudy (anonymous):

THERE IS NO X WHAT ARE YOU TALKING ABOUT

OpenStudy (spring98):

now i am so confused:(

OpenStudy (spring98):

can you just tell me how to do this? i got confused now

OpenStudy (anonymous):

I've already told you. I've explained it several ways. It's not a difficult question, all you need to do is use the distributive property and then get y alone.

OpenStudy (spring98):

? i don't see anywhere where you showed me how to get the actual y answer

OpenStudy (spring98):

can you just tell me witch one???????

OpenStudy (anonymous):

As in which answer is correct?

OpenStudy (spring98):

the steps to the correct answer!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

All you have to do is use the distributive property. Then get the y alone.

OpenStudy (spring98):

i guess i will just guess on this answer! :(